# Bandit Sequential Posted Pricing via Half-Concavity

### v1

@@ 1 Introduction
Sequential Posted Pricing (SPP) schemes are well-studied in mechanism design because of their simplicity/convenience in practice and their tractability in theory. In the basic setting, a seller wants to sell a single item to a group of buyers. The buyers arrive one-by-one and the seller presents a take-it-or-leave-it price. The first buyer with a value higher than the posted price takes the item by paying the price. The benefit of SPP is that it gives approximately-optimal revenue while being “simple”. For instance, although it is known that Myerson’s mechanism [Mye81] is optimal for selling a single item to $n$ buyers with regular distributions, this mechanism is often impractical. This is because the winner’s payment is defined via “complicated” virtual valuations and all buyer bids need to be simultaneously revealed, which is not possible in large markets. In contrast, SPP are known to give at least $1/2$-fraction of the optimal revenue for regular buyers, while having fixed prices and buyers arriving one-by-one (see the books [Rou17, Har13]). For multi-item (multi-parameters) setting, the difference between optimal and SPP mechanisms is even more stark. For instance, the optimal mechanism to sell $n$ items to a single unit-demand bidder is known to be impossible (unless $P^{NP}=P^{\#P}$) [${}^{+}$15], but we can use SPP to obtain $1/4$-fraction of the optimal revenue [CHK07, CHMS10, CMS15].
Sample Complexity. A common assumption among initial works on revenue-maximization was that the underlying distributions of the buyers are known to the seller. This is unrealistic in many applications since the distributions are unknown and need to be learnt from historical data. Inspired by this, a recent line of research studies the sample complexity for different types of mechanisms, i.e., how many samples are sufficient to learn an $\epsilon$-optimal mechanism. For instance, starting with the pioneering work in [CR14, BBHM08], several papers studied the sample complexity of Myerson’s mechanism [HMR18, DHP16, RS16, GN17], and finally [GHZ19] obtained the tight sample complexity bounds for regular buyers and for $[0,1]$ bounded-support buyers. Sample complexity of many other classes of mechanisms have been studied, e.g., [BSV18, JLX23] study second-price auctions, [BSV18, ${}^{+}$21] study posted-pricing mechanisms, [GHTZ21] study strongly-monotone auctions (which includes Myerson and SPP), and [MR16, CD17, GW21, COZ22] study the sample complexity of multi-parameter auctions.
@@
The study of learning mechanisms with bandit feedback goes back to at least [KL03], where the authors provide an $\widetilde{O}(T^{{2}/{3}})$ regret bound for learning single-buyer posted pricing mechanism. Other examples of learning auctions in this model include bandits with knapsacks [BKS18, ISSS22], reserve price for i.i.d buyers [CGM15], and non-anonymous reserve prices [${}^{+}$21].
Despite a lot of work on learning SPP mechanisms and on learning auctions with bandit feedback, learning SPP in the bandit model was unknown. In this paper we answer this question and provide near-optimal regret bounds for SPP with $n$ buyers having regular distributions or $n$ buyers having general distributions.
1.1 Model
@@
It should be noted that this is the first bandit learning algorithm for regular distributions with optimal regret bound in $T$, even for the single-buyer or i.i.d. buyers setting. A very recent paper of [LSTW23] studies “pricing query complexity” of this problem in the special case of a single-buyer. Their model is weaker than our bandit learning model since it only gives a final regret bound, i.e., one might incur a large regret during the $T$ days of learning but the final learnt prices have a low $1$-day regret. Hence, their $\Omega({1}/{\epsilon^{2}})$ lower bound for single-buyer query complexity immediately implies an $\Omega(\sqrt{T})$ lower bound for BSPP, showing tightness of the $\sqrt{T}$ factor in 1. However, their single-buyer pricing query complexity upper bounds do not apply to our stronger bandit learning model. Moreover, it’s unclear how to extend their techniques beyond a single-buyer.
Our next result gives a near-optimal regret bound for BSPP with $n$ buyers having general distributions. The authors of [LSTW23] give an $\Omega({1}/{\epsilon^{3}})$ query complexity lower bound for single-buyer setting with general distributions, implying an $\Omega(T^{2/3})$ regret lower bound for BSPP with general distributions. Our result achieves a tight upper bound dependency on $T$.
@@ 1.2 Techniques
Main Result 3 (Informal 2.4 and 3.6).
@@
Concavity is a strong property that often allows efficient learning. We show that it’s sufficient to learn a single-peaked function even if it only has half-concavity. Given a half-concave function $R(p)$ and interval $[\ell,r]$ satisfying $p^{*}\in[\ell,r]$. We first take $a=\frac{2\ell+r}{3},b=\frac{\ell+2r}{3}$ and estimate $R(a)$ and $R(b)$. If we find $R(a)$ and $R(b)$ are roughly the same, then using half-concavity we show that the values in $[b,r]$ will not be too high. Otherwise, we can distinguish between $R(a)$ and $R(b)$ and rule out one of the intervals $[\ell,a]$ or $[b,r]$ using the single-peakedness of $R(p)$; details in Section 2.
@@ 1.3 Further Related Work
Threshold Query Model. Recently, the threshold query structure that underpins the bandit feedback model of the single-buyer posted pricing problem is extensively studied. In this model, the learner queries a threshold $\tau$ and only observes $\mathbf{1}[\tau<X\sim\mathcal{D}]$. The goal is to determine the minimum number of queries to learn a key parameter(e.g., median, mean, CDF) of a distribution. For comprehensive understanding, we refer readers to [MN21, LSTW23, OGKG23] for discussions around the learning complexity of the threshold query model.
@@
A closely related problem to SPP is the Prophet Inequality (PI) problem from optimal stopping theory. In both these problems a sequence of $n$ independent buyers arrive one-by-one. However, in SPP we want to maximize the revenue and in PI we want to maximize the welfare. Interestingly, for known buyer value distributions, both these problems are equivalent [CFPV19], but this reduction requires virtual-value distributions and doesn’t work for unknown distributions. In the bandit model, PI and SPP behave very differently. For instance, [GKSW24] recently obtained $\widetilde{O}(\sqrt{T})$ regret algorithm for PI with general distributions, whereas an $\Omega(T^{2/3})$ regret lower bound exists for SPP [LSTW23]. Our $\widetilde{O}(\sqrt{T})$ regret results crucially rely on half-concavity of regular distributions.
Regular Distributions and Learning. Myerson’s regularity of distributions has been greatly studied since it was introduced in [Mye81]. In particular, it is a standard assumption in learning theory for auctions [DRY15, CR14, HMR18, DHP16, RS16, GHZ19]. For basic properties of regular distribution, and its important subclass of Monotone Hazard Rate distributions, see the books [Har13, Rou17]. A recent paper of [LSTW23], which studies pricing query complexity of the single-buyer single-item problem, proves an interesting “relative flatness” property of regular distributions in the value space. Roughly it says that if the revenue curve has nearly the same value at 4 different equidistant points then there cannot be a high revenue point in between. Comparing it to our idea of half-concavity, the two properties are in general incomparable as no one implies the other. However, we found our half-concavity to be more intuitive and convenient to work with, since proofs based on relative flatness often lead to a long case analysis.
@@ 2.1 Half-Concavity
Given both Lipschitzness and (half) concavity, the following theorem stipulates the $\widetilde{O}(\sqrt{T})$ regret bound:
@@
We will prove the three properties of half-concavity one-by-one.
@@
(iii) Half-concavity: We first provide a direct proof via the monotonicity of virtual valuation $\phi$, assuming the PDF $f(p)$ is non-zero and differentiable. An alternate prove via the concavity in the quantile space that bypasses this assumption is presented in Section A.1.
@@
    where the first inequality is from (1) and the second inequality uses $f^{\prime}(p)\leq 0$ along with the fact that $R^{\prime}(p)\geq 0$ when $p\in[0,p^{*}]$. ∎
@@ 2.3.2 Step 2: Generating New Confidence Interval
Upon executing Algorithm 2, we obtain an approximate optimal price $\hat{p}$. The subsequent stage of the algorithm uses $R(\hat{p})$ as a benchmark to establish the new upper and lower bounds of the confidence interval. Given that the function $R(p)$ is single-peaked, two independent binary search algorithms suffice to independently determine these new upper and lower bounds. Notably, the specific algorithm necessitates additional steps to manage the non-Lipschitz tail of the new confidence interval, involving numerous trivial details. Therefore, here we only present the primary lemma of Step 2, and defer the detailed algorithm and proofs to Section A.2.
@@ 3.1 Proof Overview
Step 1: Finding $\hat{p}_{i}$ with a Different Revenue Function.
Comparing to the single-buyer setting, the revenue function for sequential buyers is different. In particular, when we want to determine $\hat{p}_{i}$, the revenue function for buyer $i$ looks like
@@
The major challenge in this step is that function $R(p)$ is not always half-concave: it is half-concave in $[C,1]$, while in $[0,C]$ it is increasing but not necessary concave. This prevents us from directly applying the single-buyer algorithm. To fix this, an idea is to estimate the value of $C$ to determine the half-concave region. We show that $C$ can be estimated accurately enough in some of the cases, where then running the single-buyer algorithm is sufficient. On the other hand, in the cases where $C$ cannot be estimated, we observe an extra property of $R(p)$ that it is growing fast enough in $[0,C]$, leading to some kinds of “concavity” property in $[0,p^{*}]$. This is sufficient to show the learnability of $R(p)$. Combining the two cases completes Step 1 of our algorithm.
Step 2: Binary Search with Prices $\hat{p}_{1},\cdots,\hat{p}_{n}$.
The second step of the sub-routine is to update confidence intervals via binary searches. In this step, the main new challenge is the error from $\hat{p}_{1},\cdots\hat{p}_{n}$. For instance, while trying to update $[\ell_{i},r_{i}]$ with revenue function $R(p)=p\cdot(1-F(p))+C\cdot F(p)$, the constant $C$ comes from non-optimal prices $\hat{p}_{i+1},\cdots,\hat{p}_{n}$. Notice that $p^{*}_{i}$ is not the optimal price of $R(p)$. This creates an issue that comparing $R(\hat{p}_{i})$ with $R(p)$ for a price $p\in[\ell_{i},r_{i}]$ is not evident to show $p^{*}_{i}<p$ or $p^{*}_{i}>p$, preventing us from directly doing binary search.

### v2

@@ 1 Introduction
Sequential Posted Pricing (SPP) schemes are well-studied in mechanism design because of their simplicity/convenience in practice and their tractability in theory. In the basic setting, a seller wants to sell a single item to a group of buyers. The buyers arrive one-by-one and the seller presents a take-it-or-leave-it price. The first buyer with a value higher than the posted price takes the item by paying the price. The benefit of SPP is that it gives approximately-optimal revenue while being “simple”. For instance, although it is known that Myerson’s mechanism [Mye81] is optimal for selling a single item to $n$ buyers with regular distributions, this mechanism is often impractical. This is because the winner’s payment is defined via “complicated” virtual valuations and all buyer bids need to be simultaneously revealed, which is not possible in large markets. In contrast, SPP are known to give at least $1/2$-fraction of the optimal revenue for regular buyers, while having fixed prices and buyers arriving one-by-one (see the books [Rou17, Har13] for regular buyers, and [Yan11] for a discussion on how ironing extends this result to general buyers). For multi-item (multi-parameters) setting, the difference between optimal and SPP mechanisms is even more stark. For instance, the optimal mechanism to sell $n$ items to a single unit-demand bidder is known to be impossible (unless $P^{NP}=P^{\#P}$) [CDO+15], but we can use SPP to obtain $1/4$-fraction of the optimal revenue [CHK07, CHMS10, CMS15].
Sample Complexity. A common assumption among initial works on revenue-maximization was that the underlying distributions of the buyers are known to the seller. This is unrealistic in many applications since the distributions are unknown and need to be learnt from historical data. Inspired by this, a recent line of research studies the sample complexity for different types of mechanisms, i.e., how many samples are sufficient to learn an $\epsilon$-optimal mechanism. For instance, starting with the pioneering work in [CR14, BBHM08], several papers studied the sample complexity of Myerson’s mechanism [HMR18, DHP16, RS16, GN17], and finally [GHZ19] obtained the tight sample complexity bounds for regular buyers and for $[0,1]$ bounded-support buyers. Sample complexity of many other classes of mechanisms have been studied, e.g., [BSV18, JLX23] study second-price auctions, [BSV18, BDD+21] study posted-pricing mechanisms, [GHTZ21] study strongly-monotone auctions (which includes Myerson and SPP), and [MR16, CD17, GW21, COZ22] study the sample complexity of multi-parameter auctions.
@@
The study of learning mechanisms with bandit feedback goes back to at least [KL03], where the authors provide an $\widetilde{O}(T^{{2}/{3}})$ regret bound for learning single-buyer posted pricing mechanism. Other examples of learning auctions in this model include bandits with knapsacks [BKS18, ISSS22], reserve price for i.i.d buyers [CGM15], and non-anonymous reserve prices [NGW+21].
Despite a lot of work on learning SPP mechanisms and on learning auctions with bandit feedback, learning SPP in the bandit model was unknown. We answer this question and provide near-optimal regret bounds for SPP with $n$ buyers having regular or general distributions.
1.1 Model and Results
@@
It should be noted that this is the first bandit learning algorithm for regular distributions with optimal regret bound in $T$, even for the single-buyer or i.i.d. buyers setting. A very recent paper of [LSTW23b] studies “pricing query complexity” of this problem in the special case of a single-buyer. Their model is weaker than our bandit learning model since it only gives a final regret bound, i.e., one might incur a large regret during the $T$ days of learning but the final learnt prices have a low $1$-day regret. Hence, their $\Omega({1}/{\epsilon^{2}})$ lower bound for single-buyer query complexity immediately implies an $\Omega(\sqrt{T})$ lower bound for BSPP, showing tightness of the $\sqrt{T}$ factor in 1. However, their single-buyer pricing query complexity upper bounds do not apply to our stronger bandit learning model. Moreover, it’s unclear how to extend their techniques beyond a single-buyer.
Our next result gives a near-optimal regret bound for BSPP with $n$ buyers having general distributions. The authors of [LSTW23b] give an $\Omega({1}/{\epsilon^{3}})$ query complexity lower bound for single-buyer setting with general distributions, implying an $\Omega(T^{2/3})$ regret lower bound for BSPP with general distributions. Our result achieves a tight upper bound dependency on $T$.
@@ 1.2 Techniques
Main Result 3 (Informal 2.4 and 3.9).
@@
Concavity is a strong property that often allows efficient learning. We show that it’s sufficient to learn a single-peaked function even if it is only half-concave. The high-level intuition for learning a half-concave follows from the standard recursive algorithm for learning concave functions: first, consider a (fully) concave function $R(x)$ defined on interval $[\ell,r]$. We set $a=\frac{2\ell+r}{3}$ and $b=\frac{\ell+2r}{3}$, and test $R(a)$ and $R(b)$ with sufficiently many samples. There can be three cases:
  • Case 1: $R(a)<R(b)$. We can drop $[\ell,a]$ and recurse on $[a,r]$.
  • Case 2: $R(a)>R(b)$. We can drop $[b,r]$ and recurse on $[\ell,b]$.
  • Case 3: $R(a)\approx R(b)$. Concavity implies that $R(x)$ is nearly a constant for all $x\in[\ell,r]$, so we are done.
Now suppose $R(x)$ is only half-concave. We can still perform a similar algorithm: the first two cases use single-peakedness, so they remain the same. For the third case, half-concavity still guarantees that $R(x)$ can’t be too high for $x\in[b,r]$, so we can drop $[b,r]$ and recurse on $[\ell,b]$.
Our final algorithm and proofs combine the above Case 2 and 3; see details in Section 2.
@@ 1.3 Further Related Work
Threshold Query Model. Recently, the threshold query structure that underpins the bandit feedback model of the single-buyer posted pricing problem is extensively studied. In this model, the learner queries a threshold $\tau$ and only observes $\mathbf{1}[\tau<X\sim\mathcal{D}]$. The goal is to determine the minimum number of queries to learn a key parameter (e.g., median, mean, or CDF) of a distribution. We refer the readers to [MN21, LSTW23b, OGKG23, LSTW23a] for learning complexity of the threshold query model.
@@
A closely related problem to SPP is the Prophet Inequality (PI) problem from optimal stopping theory. In both these problems a sequence of $n$ independent buyers arrive one-by-one. However, in SPP we want to maximize the revenue and in PI we want to maximize the welfare. Interestingly, for known buyer value distributions, both these problems are equivalent [CFPV19], but this reduction requires virtual-value distributions and doesn’t work for unknown distributions. In the bandit model, PI and SPP behave very differently. For instance, [GKSW24] recently obtained $\widetilde{O}(\sqrt{T})$ regret algorithm for PI with general distributions, whereas an $\Omega(T^{2/3})$ regret lower bound exists for SPP [LSTW23b]. Our $\widetilde{O}(\sqrt{T})$ regret results crucially rely on half-concavity of regular distributions.
Regular Distributions and Learning. Myerson’s regularity of distributions has been greatly studied since it was introduced in [Mye81]. In particular, it is a standard assumption in learning theory for auctions [DRY15, CR14, HMR18, DHP16, RS16, GHZ19]. For basic properties of regular distribution, and its important subclass of Monotone Hazard Rate distributions, see the books [Har13, Rou17]. A recent paper of [LSTW23b], which studies pricing query complexity of the single-buyer single-item problem, proves an interesting “relative flatness” property of regular distributions in the value space. Roughly it says that if the revenue curve has nearly the same value at 4 different equidistant points then there cannot be a high revenue point in between. Comparing it to our idea of half-concavity, the two properties are in general incomparable as no one implies the other. However, we found our half-concavity to be more intuitive and convenient to work with, since proofs based on relative flatness often lead to a long case analysis.
@@ 2.1 Half-Concavity
Given both Lipschitzness and (half) concavity, the following theorem gives $\widetilde{O}(\sqrt{T})$ regret.
@@
We first provide a proof assuming the PDF $f(p)$ is non-zero and differentiable. We will prove the three properties of half-concavity one-by-one.
@@
(iii) Half-concavity: We first provide a direct proof via the monotonicity of virtual valuation $\phi$.
@@
    where the first inequality is from (1) and the second inequality uses $f^{\prime}(p)\leq 0$ along with the fact that $R^{\prime}(p)\geq 0$ when $p\in[0,p^{*}]$.
The above proof requires the PDF $f(p)$ to be non-zero and differentiable. Below we provide an alternate proof via the concavity in the quantile space, which bypasses this assumption for $f(p)$. We keep both the proofs here since the reader may find one proof more instructive than the other.
An alternate proof for half-concavity. We prove half-concavity by contradiction. Assume there exist $0\leq a<b<c\leq p^{*}$ satisfying $c=b+t(b-a)$ and $R(c)>R(b)+t\cdot\big{(}R(b)-R(a)\big{)}$.
$R(c)$ $~{}>~{}R(b)+t\cdot(R(b)-R(a))$
$~{}=~{}b\cdot q(b)+t\cdot\big{(}b\cdot q(b)-a\cdot q(a)\big{)}$
$~{}=~{}\big{(}(t+1)b-ta\big{)}\cdot\big{(}(t+1)q(b)-tq(a)\big{)}$
$~{}=~{}c\cdot\overline{q}~{}\geq~{}c\cdot q(c)~{}=~{}R(c).\qed$
@@ 2.3.2 Step 2: Generating New Confidence Interval
Upon executing Algorithm 2, we obtain an approximate optimal price $\hat{p}$. The subsequent stage of the algorithm uses $R(\hat{p})$ as a benchmark to establish the new upper and lower bounds of the confidence interval. Given that the function $R(p)$ is single-peaked, two independent binary search algorithms suffice to independently determine these new upper and lower bounds, leading to the following 2.9:
@@
It only remains to prove 2.9. We first give the following Algorithm 3, and then show Algorithm 3 is the desired algorithm for 2.9.
1 Let $\delta=\frac{\epsilon}{100}$ be the scaled error parameter.
// First binary search to determine $\ell^{\prime}$
3 Let $\ell_{b1}\leftarrow\ell,r_{b1}\leftarrow\hat{p}$.
4 while $r_{b1}-\ell_{b1}\geq{T^{-100}}$ do
5       Let $m=\frac{\ell_{b1}+r_{b1}}{2}$.
8
9Let $\ell^{\prime}\leftarrow\ell_{b1}$
// Second binary search to determine $r^{\prime}$
11 if $\hat{R}(r-{T^{-100}})\geq\hat{R}(\hat{p})-2\delta$ then
       // Special Case: original upper bound is good
12       Let $r^{\prime}\leftarrow r$
13else
14       Let $\ell_{b2}\leftarrow\hat{p},r_{b2}\leftarrow r-{T^{-100}}$.
15       while $r_{b2}-\ell_{b2}\geq{T^{-100}}$ do
16             Let $m=\frac{\ell_{b2}+r_{b2}}{2}$.
19
20      Let $r^{\prime}\leftarrow r_{b2}$.
Output: $[\ell^{\prime},r^{\prime}]$.
ALGORITHM 3 Getting $[\ell^{\prime},r^{\prime}]$ via Binary Search
We first give the following concentration bound for Algorithm 3:
We first give the following claim:
Claim 2.10.
Proof.
Now we prove 2.9 assuming 2.10 holds.
Proof of 2.9.
Next, we show $p^{*}\in[\ell^{\prime},r^{\prime}]$. We first prove $p^{*}\geq\ell^{\prime}$. In Algorithm 3, the only way we update $\ell^{\prime}$ is that we observe $\hat{R}(m)<\hat{R}(\hat{p})-2\delta$ in the first binary search, and $\ell^{\prime}$ is updated to at least $m$. This is feasible because 2.10 gives $R(m)<R(\hat{p})\leq R(p^{*})$, and the single-peakedness of $R(p)$ guarantees $m\leq p^{*}$. The proof of $p^{*}\leq r^{\prime}$ is symmetric.
We first prove that $R(\hat{p})-R(\ell^{\prime})\leq 5\delta$. The binary search subroutine guarantees that $\hat{R}(r_{b1})\geq\hat{R}(\hat{p})-2\delta$ after the binary search loop is finished. Then, 2.10 ensures that $R(\hat{p})-R(r_{b1})\leq 4\delta$. Since $\ell^{\prime}$ is set to be $\ell_{b1}>r_{b1}-{T^{-100}}$, the 1-Lipschitzness of $R(p)$ in $[\ell_{b1},\hat{p}]$ guarantees that
$R(r^{\prime}-{T^{-100}})~{}\geq~{}R(\ell_{b2})~{}\geq~{}R(\hat{p})-4\delta.$
Otherwise, the 1-Lipschitzness gives
@@ 3.1 Proof Overview
Step 1: Finding $\hat{p}_{i}$ with a Different Revenue Function. Comparing to the single-buyer setting, the revenue function for sequential buyers is different. In particular, when we want to determine $\hat{p}_{i}$, the revenue function for buyer $i$ looks like
@@
The major challenge in this step is that function $R(p)$ is not always half-concave: it is half-concave in $[C,1]$, while in $[0,C]$ it is increasing but not necessary concave. This prevents us from directly applying the single-buyer algorithm. To fix this, an idea is to estimate the value of $C$ to determine the half-concave region. We show that $C$ can be estimated accurately enough in some of the cases, where then running the single-buyer algorithm is sufficient. On the other hand, in the cases where $C$ cannot be estimated, we observe an extra property of $R(p)$ that it is growing fast enough in $[0,C]$, leading to a “generalized half-concavity” property in $[0,p^{*}]$. This is sufficient to show the learnability of $R(p)$. Combining the two cases completes Step 1 of our algorithm.
Step 2: Binary Search with Prices $\hat{p}_{1},\cdots,\hat{p}_{n}$. The second step of the sub-routine is to update confidence intervals via binary searches. In this step, the main new challenge is the error from $\hat{p}_{1},\cdots\hat{p}_{n}$. For instance, while trying to update $[\ell_{i},r_{i}]$ with revenue function $R(p)=p\cdot(1-F(p))+C\cdot F(p)$, the constant $C$ comes from non-optimal prices $\hat{p}_{i+1},\cdots,\hat{p}_{n}$. Notice that $p^{*}_{i}$ is not the optimal price of $R(p)$. This creates an issue that comparing $R(\hat{p}_{i})$ with $R(p)$ for a price $p\in[\ell_{i},r_{i}]$ is not evident to show $p^{*}_{i}<p$ or $p^{*}_{i}>p$, preventing us from directly doing binary search.
@@
3.2 Notation
  • $\mathsf{Loss}(S_{i})$: We define $\mathsf{Loss}(S_{i}):=R(1,\cdots,1,p^{*}_{i},\cdots,p^{*}_{n})-R(1,\cdots,1,\hat{p}_{i},\cdots,\hat{p}n)$, i.e., $\mathsf{Loss}(S_{i})$ represents the loss from $S_{i}$ when playing prices $\hat{p}_{i},\cdots,\hat{p}_{n}$.
    The intuition of function $R_{i}(p)$ is: When finding $\hat{p}_{i}$, naturally we hope to fix other prices to avoid interference from other buyers. For a buyer $j$ before $i$, we set $p_{j}$ to be $r^{s}_{j}$ to maximize the probability of testing $i$. For buyer $k$ from suffix $S_{i+1}$, we set $p_{k}$ to be the settled $\hat{p}_{k}$ to minimize the error from the suffix.
    The intuition of $R^{*}_{i}(p)$ is: notice that $\arg\max_{p}r^{*}_{i}(p)=p^{*}_{i}$. A key idea of our algorithm is to approach $p^{*}_{i}$ via estimating $R^{*}_{i}(p)$. It should be noted that $R^{*}_{i}(p)$ can be also written as $R^{*}_{i}(p)=R_{i}(p)+P_{i}\cdot F_{i}(p)\cdot\mathsf{Loss}(S_{i+1}).$ We will mainly use the second definition in our proofs.